# Hyperband-based Bayesian Optimization for Black-box Prompt Selection

### v1

@@ Appendix C Multi-Fidelity over Validation Instances
In this section, we discuss how one can define a multi-fidelity schedule for prompt selection where the fidelity is the number of validation instances. Let $\mathcal{D}_{\mathrm{valid}}=\{(x_{i},y_{i})\}_{i=1}^{n_{\mathrm{valid}}}$ denote the validation set containing $n_{\mathrm{valid}}$ input-output instances in natural language on which the LLM $h_{p}$ configured to use a given prompt $p\in\mathcal{P}$ is evaluated. Recall our goal of identifying the optimal prompt: $\operatorname*{arg\,min}_{p\in\mathcal{P}}\mathbb{E}_{(x,y)\sim\mathbb{P}_{xy}}\left[l(y,h_{p}(x))\right]$. Here, the expectation is taken over all input output instances from a data generating distribution $\mathbb{P}_{xy}$ and $l$ is the point-wise loss function used to compare the LLMs output $h_{p}(x)$ to the ground truth $y$.
@@
The idea of multi-fidelity techniques is to speed up the evaluation of prompts by using fewer validation instances during evaluation. Let $\mathcal{V}=\{1,\ldots,n_{\mathrm{valid}}\}$ denote the index set corresponding to the indices of validation instances. A simple way to speed up evaluation is to use a random subset, $\mathcal{S}\subset\mathcal{V}$ of validation instances: $\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}l(y_{i},h_{p}(x_{i}))$. This was for example done in the evaluation protocol of the benchmark study in Wu et al. (2024). This has the following downsides: (1) It is a priori not clear how many instances are needed to obtain a precise estimate of the validation performance of prompts as this depends on the concrete LLM, downstream task, and variation of the true performance over prompts. (2) Using a fixed sub-sample is inefficient because all prompts are evaluated on the same number of validation instances. However, poor performing prompts can be identified by using few validation instances, whereas well performing prompts should be evaluated on more validation instances to be able to differentiate between them and not risk a sub-optimal selection.
@@
Shi et al. (2024) made the connection between multi-fidelity prompt selection and best arm identification from the multi-armed bandit literature. In this setting, “pulling an arm” refers to evaluating a prompt on a validation instance. The goal is to identify the best arm within a limited budget of evaluations. A well-known algorithm from the bandit literature is given by Successive Halving (SH, Karnin et al. 2013; Jamieson & Talwalkar 2016). In our setting of prompt selection, the idea of SH is to efficiently identify the best performing prompt (arm) under a given budget constraint of total LLM calls (pulls). Given a total budget $B$ of LLM calls and $n\coloneqq|\mathcal{P}|$ prompts, SH starts by allocating a budget of $b\coloneqq B/(n\log_{2}(n))$ LLM calls to each prompt. After having evaluated each prompt on $b$ instances, the lower half of bad performing prompts are discarded and the process repeats, doubling the number of LLM calls for the remaining prompts in the next stage. This process in general repeats until a single prompt remains.
SH can be employed both in a stochastic and non-stochastic setting. The stochastic setting (Karnin et al., 2013) is formally characterized by the following: (1) Losses are i.i.d. samples from a probability distribution. (2) Each arm has a fixed expected loss $\mu_{p}$. (3) The goal is to identify the arm with the lowest expected loss.
@@
Jamieson & Talwalkar (2016) introduced SH in the non-stochastic setting when applying it to the problem of Hyperparameter Optimziation, where the budget can, for example, be the size of the training set used to train an algorithm or the number of epochs used to train a neural network. This non-stochastic setting formally is characterized by the following: (1) Losses are real numbers chosen by an oblivious adversary. (2) Each arm has a limit $\nu_{p}$ of its loss sequence as the number of evaluations go to infinity. (3) The goal is to identify the arm with the lowest loss limit.
@@
Within the non-stochastic setting, Li et al. (2018) introduced Hyperband (HB) albeit it is in general also applicable to the stochastic setting. Recall that SH requires the overall budget $B$ and the number of prompts $n$ as input which determines the starting budget $b$ (number of validation instances) prompts are evaluated on. However, for a given overall budget $B$ it is not clear whether one should evaluate may prompts using on average few validation instances (allowing for better exploration of the search space at the risk of noisy validation error estimates) or whether one should consider a smaller number of prompts using on average more validation instances (focusing on fewer prompts but obtaining more precise validation error estimates).
@@
Our proposal to use HB as a multi-fidelity scheduler for prompt selection improves over SH by addressing this trade-off between the number of prompts to explore and the amount of resources to allocate to each configuration. It does this by running multiple SH brackets, each with a different initial number of prompts and per-prompt budget. This allows HB to efficiently explore the search space, quickly discarding poor prompts while allocating more resources to promising ones while hedging against a poor choice of the number of prompts and per-prompt budget.
@@
We presented pseudo code for HB adapted to the problem of prompt selection in Algorithm 1 and in Table 4 we present an exemplary schedule describing how the number of prompts used in each stage of each bracket of the algorithm relates to the number of validation instances used to evaluate the prompts. Regarding the output of the algorithm we note that there is a critical design decision: The vanilla HB (Li et al., 2018) algorithm would output the configuration with the smallest validation error seen so far. However, in the context of prompt selection, the validation errors are determined based on different numbers of validation instances varying between stages of brackets ($\{s_{\max},s_{\max}-1,\ldots,0\}$). To allow for a robust selection of the optimal prompt, we always return the prompt with the lowest validation error among all prompts that were evaluated on the full validation set111or evaluated on the largest subset used so far when assessing the anytime performance of the algorithm. This is crucial for robust performance of HB in the context of prompt selection which we ablate in Appendix E.4.
@@
Table 4: Exemplary HB schedule for prompt selection assuming a minimum budget of $b_{\min}=10$ validation instances, a maximum number of $n_{\mathrm{valid}}=80$ validation instances being available in total and a halving parameter of $\eta=2.0$.
Bracket ($s$) Stage ($i$) #Instances ($b$) #Prompts ($n$) 3 0 10 8 3 1 20 4 3 2 40 2 3 3 80 1 2 0 20 6 2 1 40 3 2 2 80 1 1 0 40 4 1 1 80 2 0 0 80 4
@@
Our modified HB has three inputs: $n_{\mathrm{valid}}$ the total number of validation instances available (depending on the task),$b_{\mathrm{min}}$ a lower bound to the number of validation instances used to evaluate a prompt, $\eta$ the halving parameter of the SH subroutine. HB makes use of three functions:
@@
Moreover, when moving from one stage to another stage, there are two possibilities how to construct the subset of validation instances used in the next stage: (1) Simply draw a sample. (2) Keep the already used validation instances of the previous stage and only sample the remaining number of additional validation instances needed to fill the current stage from the remaining yet not used validation instances. The second option is highly desirable to reduce the total number of LLM calls if we cache the evaluation of prompts. How reasonable this is of course depends on the stochasticity of the output of the LLM for a given prompt on a given instance (which depends on the temperature). However, if we assume deterministic outcomes, we can reduce the total number of LLM calls used in HB drastically (i.e., roughly by a factor of $\eta$), In HbBoPs we cache the output for a given prompt and validation instance and reuse the already used validation instances of the previous stage and only sample the remaining additionally needed validation instances. We ablate this design decision in Appendix E.4.
@@ D.1 Tasks
Table 5 reports characteristics on the benchmark tasks used by us. We use the official train, validation and test splits from AI2’s Reasoning Challenge. However, during inspection we noticed that for some reason, AI2’s Reasoning Challenge includes very few instances with choices named “1”, “2”, “3”, “4” instead of “A”, “B”, “C”, “D” and very few instances with five choices instead of four which we excluded from the splits for consistency. GSM8K officially only contains a train and test split. We therefore sampled $1319$ instances from the train split uniformly at random as a separate validation split to have a similar sized validation split as the test split. Regarding all other tasks from the BBH and instruction induction benchmark we use the splits as processed and used by Wu et al. (2024). Note that Wu et al. (2024), however, further sub-sampled their validation split to only use $20$ validation instances in their experiments. For each task, the training split was used to generate the instructions via APE’s (Zhou et al., 2022) forward mode and to select instances for few-shot exemplars, whereas the validation split was used during optimization and the test split was used to assess unbiased performance.
Table 5: Characteristics of Tasks used in the Experiments
Task Setting $n_{\mathrm{train}}$ $n_{\mathrm{valid}}$ $n_{\mathrm{test}}$ AI2 ARC multiple choice question answering $1094$ $291$ $1144$ GSM8k grade school math questions $6154$ $1319$ $1319$ antonyms find antonym of word $2073$ $519$ $100$ larger animal select larger of two animals $2422$ $606$ $100$ negation negate a sentence $723$ $181$ $100$ object counting count number of objects $560$ $140$ $100$ orthography starts with output all words starting with a given letter $2400$ $600$ $100$ second word letter output the second letter of a word $2644$ $662$ $100$ sentiment sentiment analysis of movie rating $933$ $234$ $100$ word unscrambling build a word from scrambled letters $5627$ $1407$ $100$
@@ D.2 LLMS
We use Claude 3 Haiku (Anthropic, 2024), LLAMA3 8B Instruct (Dubey et al., 2024), and Mistral 7B Instruct (Jiang et al., 2023) with the following default hyperparameters (Claude 3 Haiku: max tokens $=200$, temperature $=0.5$, top p $=1.0$, top $k=250$; LLAMA3 8B Instruct: max tokens $=512$, temperature $=0.5$, top $p=0.9$; Mistral 7B Instruct: max tokens $=512$, temperature $=0.5$, top p $=0.9$, top $k=50$). For GSM8K we increase max tokens to $1024$
@@ D.3 Methods
We run all methods as described in Section 4. All full-fidelity BO methods are implemented within BoTorch (Balandat et al., 2020). We run HDBO (Hvarfner et al., 2024) to have a simple yet well-performing ”high-dimensional” BO baseline. Hvarfner et al. (2024) recently challenged the general belief that vanilla BO does not perform well for high-dimensional functions by training a GP with priors over kernel and likelihood parameters adjusted to reflect the dimensionality of the problem which resulted in strong BO performance on high-dimensional functions. We run EASE as implemented in their official code base [^2] https://github.com/ZhaoxuanWu/EASE-Prompt-Optimization/blob/e3514de58bd682ebc5ea46fe890481f2b92e5589/experiments/LlamaForMLPRegression.py, MIPROv2 as implemented in DSPy [^3] https://github.com/stanfordnlp/dspy/blob/425b6f07d5cf0530f5a5566ad4f247b15aecb522/dspy/teleprompt/mipro_optimizer_v2.py and TRIPLE-GSE as implemented in their official code base [^4] https://github.com/ShenGroup/TRIPLE/blob/06264a97b4dd766c9a88afc24058627fac0f223d/src/bandit/contextual/gse.py. Regarding TRIPLE-GSE we noticed that Shi et al. (2024) did not describe in detail the model they are using to predict the expected prompt performance (in the main paper, they state that one can use a linear model or MLP). In their implementation, however, they use an ensemble of a Bayesian ridge regression model, gradient boosting regression model and an MLP where weights for the ensemble are determined based on each model’s $R^{2}$ performance on a validation set. We did not change this modeling approach when running TRIPLE-GSE. We implement our HbBoPs in GPytorch (Gardner et al., 2018) and run it as described in Section 4. Moreover, to hedge against poor model based proposals early on, we perform random interleaving as described in Falkner et al. (2018) for each proposal with a probability of $\rho=0.1$.
@@ E.1 Main Results
Here we provide additional analyses of the main results reported in Section 5.1. To test whether HbBoPs outperforms all other methods with respect to validation and test performance for different fractions of budget, we conduct a linear mixed effects model analysis. We model the unaggregated performance as a function involving random intercepts for each benchmark scenario (benchmark task and LLM combination). This approach is sensible as each method has been run repeatedly on the same benchmark scenario with different random seeds that effect for example initial designs. To test the global hypothesis that there is an effect of the method on performance, we test an intercept model against a model including an effect of the factor method. If we reject the null hypothesis we proceed with a Tukey post hoc test (corrected for multiple testing) to test each method against HbBoPs. We test at the $\alpha=0.01$ level.
For the validation performance at a fraction of $1.00$, we reject the global null hypothesis of no effect of methods ($\chi^{2}(8)=672.95,p<1\mathrm{e}{-4}$). The pairwise results are:
@@
We can see that HbBoPs outperforms all methods significantly with the exception of TRIPLE-SE. While the effect is positive, i.e., HbBoPs improves over TRIPLE-SH, the effect is not strong enough to be considered statistically significant at the $\alpha=0.01$ level. For brevity, we do not include results for fractions of $0.25$ or $0.50$ here, where we observed similar results but HbBoPs more strongly outperforming the other methods.
Table 6: Normalized validation error of each method on each benchmark for Claude 3 Haiku. Averaged over repetitions. Standard errors in parentheses. For visual analysis, we highlight all methods that have a mean performance that is less or equal to the mean performance of the best method plus two times its standard error.
@@
Table 7: Normalized validation error of each method on each benchmark for LLAMA3 8B Instruct. Averaged over repetitions. Standard errors in parentheses. For visual analysis, we highlight all methods that have a mean performance that is less or equal to the mean performance of the best method plus two times its standard error.
@@
Table 8: Normalized validation error of each method on each benchmark for Mistral 7B Instruct. Averaged over repetitions. Standard errors in parentheses. For visual analysis, we highlight all methods that have a mean performance that is less or equal to the mean performance of the best method plus two times its standard error.
Benchmark AI2 ARC GSM8K antonyms larger animal negation object counting orthography starts with second word letter sentiment word unscrambling RS 0.329 (0.035) 0.160 (0.021) 0.020 (0.002) 0.043 (0.006) 0.006 (0.003) 0.040 (0.004) 0.164 (0.018) 0.171 (0.016) 0.098 (0.013) 0.224 (0.017) vanilla BO 0.359 (0.038) 0.166 (0.021) 0.017 (0.002) 0.040 (0.008) 0.008 (0.003) 0.033 (0.005) 0.117 (0.020) 0.136 (0.013) 0.085 (0.016) 0.161 (0.019) HDBO 0.217 (0.045) 0.105 (0.019) 0.013 (0.002) 0.021 (0.005) 0.009 (0.003) 0.028 (0.005) 0.120 (0.018) 0.085 (0.017) 0.042 (0.013) 0.108 (0.016) BOPCA 0.246 (0.044) 0.108 (0.019) 0.015 (0.002) 0.021 (0.006) 0.011 (0.004) 0.032 (0.006) 0.112 (0.021) 0.125 (0.019) 0.048 (0.013) 0.136 (0.021) EASE 0.384 (0.033) 0.133 (0.014) 0.022 (0.002) 0.053 (0.008) -0.000 (0.000) 0.035 (0.005) 0.147 (0.023) 0.183 (0.014) 0.110 (0.014) 0.095 (0.021) MIPROv2 0.323 (0.041) 0.122 (0.017) 0.013 (0.002) 0.032 (0.005) 0.011 (0.004) 0.032 (0.005) 0.125 (0.014) 0.158 (0.016) 0.068 (0.014) 0.163 (0.017) TRIPLE-SH 0.024 (0.011) 0.008 (0.003) 0.041 (0.005) 0.012 (0.003) 0.044 (0.008) 0.057 (0.006) 0.005 (0.004) 0.044 (0.012) 0.055 (0.010) 0.103 (0.018) TRIPLE-GSE 0.038 (0.016) 0.001 (0.000) 0.042 (0.005) 0.022 (0.006) 0.055 (0.010) 0.067 (0.010) 0.002 (0.000) 0.072 (0.015) 0.084 (0.017) 0.145 (0.022) HbBoPs 0.107 (0.036) 0.000 (0.000) 0.011 (0.001) 0.012 (0.003) 0.012 (0.005) 0.025 (0.005) 0.024 (0.010) 0.067 (0.013) 0.030 (0.011) 0.061 (0.014)
@@
We perform the same analysis for the test performance at a fraction of $1.00$ and reject the global null hypothesis of no effect of methods ($\chi^{2}(8)=288.36,p<1\mathrm{e}{-4}$). The pairwise results are:

### v2

@@ Appendix C Multi-Fidelity over Validation Instances
In this section, we discuss how one can define a multi-fidelity schedule for prompt selection, in which the fidelity parameter is the number of validation instances. Let $\mathcal{D}_{\mathrm{valid}}=\{(x_{i},y_{i})\}_{i=1}^{n_{\mathrm{valid}}}$ denote the validation set containing $n_{\mathrm{valid}}$ input-output instances in natural language on which the LLM $h_{p}$ configured to use a given prompt $p\in\mathcal{P}$ is evaluated. Recall our goal of identifying the optimal prompt: $\operatorname*{arg\,min}_{p\in\mathcal{P}}\mathbb{E}_{(x,y)\sim\mathbb{P}_{xy}}\left[l(y,h_{p}(x))\right]$. Here, the expectation is taken over all input output instances from a data-generating distribution $\mathbb{P}_{xy}$ and $l$ is the pointwise loss function used to compare the LLM’s output $h_{p}(x)$ to the ground truth $y$. We want to identify the best performing prompt while minimizing the number of LLM calls for evaluation, given the significant costs in both time, but especially query expenses associated with LLM black-box APIs.
@@
The idea of multi-fidelity techniques is to speed up and reduce the cost of the evaluation of prompts by using fewer validation instances during evaluation. Let $\mathcal{V}=\{1,\ldots,n_{\mathrm{valid}}\}$ denote the index set corresponding to the indices of validation instances. A simple way to reduce cost of evaluation is to use a random subset, $\mathcal{S}\subset\mathcal{V}$ of validation instances: $\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}l(y_{i},h_{p}(x_{i}))$. This was for example done in the evaluation protocol of the benchmark study in Wu et al. (2024). However, it comes with the following downsides: (1) It is a priori not clear how many instances are needed to obtain an accurate estimate of the validation error of prompts as this depends on the concrete LLM, downstream task, and variation of the true performance over prompts. (2) Using a fixed (sub-)sample is inefficient because all prompts are evaluated on the same number of validation instances. However, poor-performing prompts can be identified using few validation instances, whereas well-performing prompts should be evaluated on many validation instances to differentiate between them and not risk suboptimal selection.
@@
Shi et al. (2024) made the connection between multi-fidelity prompt selection and best arm identification from the multi-armed bandit literature. In this setting, “pulling an arm” refers to evaluating a prompt on a validation instance. The goal is to identify the best arm within a limited budget of evaluations. A well-known algorithm from the bandit literature is given by Successive Halving (SH, Karnin et al. 2013; Jamieson & Talwalkar 2016). In our setting of prompt selection, the idea of SH is to efficiently identify the best performing prompt (arm) under a given budget constraint of total LLM calls (pulls). Given a total budget $B$ of LLM calls and $n\coloneqq|\mathcal{P}|$ prompts, SH starts by allocating a budget of $b\coloneqq B/(n\log_{2}(n))$ LLM calls to each prompt. After having evaluated each prompt on $b$ instances, the lower half of bad performing prompts are discarded and the process repeats, doubling the number of LLM calls for the remaining prompts in the next stage. This process in general repeats until a single prompt remains. SH can be employed in both a stochastic and non-stochastic setting. The stochastic setting (Karnin et al., 2013) is formally characterized by the following: (1) Losses are i.i.d. samples from a probability distribution. (2) Each arm has a fixed expected loss $\mu_{p}$. (3) The goal is to identify the arm with the lowest expected loss.
@@
Jamieson & Talwalkar (2016) introduced SH in the non-stochastic setting when applying it to the problem of hyperparameter optimization, where the budget can, for example, be the size of the training set used to train an algorithm or the number of epochs used to train a neural network. This non-stochastic setting formally is characterized by the following: (1) Losses are real numbers chosen by an oblivious adversary. (2) Each arm has a limit $\nu_{p}$ of its loss sequence as the number of evaluations go to infinity. (3) The goal is to identify the arm with the lowest loss limit.
@@
Within the non-stochastic setting, Li et al. (2018) introduced Hyperband (HB) although it is in general also applicable to the stochastic setting. Recall that SH requires the overall budget $B$ and the number of prompts $n$ as input parameters. These determine the starting budget $b$ (number of validation instances) prompts are evaluated on. However, for a given overall budget $B$ it is a priori not clear whether one should evaluate many prompts using on average few validation instances (resulting in better exploration of the search space at the risk of noisy validation error estimates) or whether one should consider a smaller number of prompts using on average more validation instances (focusing on fewer prompts but obtaining more accurate validation error estimates).
@@
Our proposal to use HB as a multi-fidelity scheduler for prompt selection improves over SH by addressing this trade-off between the number of prompts to explore and the amount of resources to allocate to each configuration. It does so by running multiple SH brackets, each with a different initial number of prompts and initial per-prompt budget. This allows HB to efficiently explore the search space, quickly discarding poor prompts while allocating more resources to promising ones, therefore hedging against a poor choice of the number of prompts and per-prompt budget.
@@
We present pseudocode for HB adapted to the problem of black-box prompt selection in Algorithm 1 (as explained in Section 3.4, HbBoPs replaces the random proposal mechanism of HB by a BO proposal) and in Table 4, we present an exemplary schedule describing how the number of prompts used in each stage of each bracket of the algorithm relates to the number of validation instances used to evaluate the prompts. Regarding the output of the algorithm, we note that there is a critical design decision: The vanilla HB (Li et al., 2018) algorithm would output the configuration with the smallest validation error observed so far. However, in the context of prompt selection, validation errors are determined based on different numbers of validation instances varying between stages of brackets ($\{s_{\max},s_{\max}-1,\ldots,0\}$). To allow for a robust selection of the optimal prompt, we always return the prompt with the lowest validation error among all prompts that were evaluated on the full validation set [^2] Or evaluated on the largest subset used so far when assessing the anytime performance of the algorithm.. This is crucial for robust performance of HB in the context of prompt selection, which we demonstrate in Appendix E.4.
@@
Table 4: Exemplary HB schedule for black-box prompt selection assuming a minimum budget of $b_{\min}=10$ validation instances, a maximum number of $n_{\mathrm{valid}}=80$ validation instances being available in total, and a halving parameter of $\eta=2.0$.
Bracket ($s$) Stage ($i$) #Instances ($b$) #Prompts ($n$)
3 0 10 8
3 1 20 4
3 2 40 2
3 3 80 1
2 0 20 6
2 1 40 3
2 2 80 1
1 0 40 4
1 1 80 2
0 0 80 4
@@
Our HB adapted to black-box prompt selection has three inputs: $n_{\mathrm{valid}}$ the total number of validation instances available (depending on the task), $b_{\mathrm{min}}$ a lower bound on the number of validation instances used for prompt evaluation, and $\eta$ the halving parameter of the SH subroutine. HB makes use of three functions:
@@
Moreover, when moving from one stage to another stage, there are two possibilities how to construct the subset of validation instances used in the next stage: (1) Simply draw a sample. (2) Keep the already used validation instances of the previous stage and only sample the remaining number of additional validation instances needed to fill the current stage from the remaining yet not used validation instances. The second option is highly desirable to reduce the total number of LLM calls if we cache the evaluation of prompts. The validity of this modeling choice naturally depends on the degree of stochasticity in the LLM’s output for a given prompt-instance pair, which in turn is influenced by the sampling temperature. However, if we assume reasonably deterministic outcomes, we can reduce the total number of LLM calls used in HB drastically (i.e., roughly by a factor of $\eta$), In HbBoPs we cache the output for a given prompt and validation instance and reuse the already used validation instances of the previous stage and only sample the remaining needed validation instances. We ablate this design decision in Appendix E.4.
@@ D.1 Tasks
Table 5 reports characteristics on the benchmark tasks used in our experiments. For AI2 ARC, we use the official train, validation and test splits from AI2’s Reasoning Challenge. However, during inspection we noticed that for some reason, AI2’s Reasoning Challenge includes a few instances with choices named “1”, “2”, “3”, “4” instead of “A”, “B”, “C”, “D” and a few instances with five choices instead of four which we excluded from the splits for consistency. GSM8K officially only contains a train and test split. We sampled $1319$ instances from the train split uniformly at random to create a validation set of comparable size to the test set. For all other tasks from the BBII subset of the BIG-bench and instruction induction benchmarks, we use the splits as proposed by Wu et al. (2024). Unlike Wu et al. (2024), who used only 20 validation instances by further sub-sampling the validation splits, we retain larger validation splits to reduce noise and improve the reliability of performance estimates. For each task, the training split was used to generate instructions for the prompts via APE’s (Zhou et al., 2023) forward mode and to select instances for few-shot exemplars. Instructions are generated using APE’s forward mode, where Claude 3 Sonnet (Anthropic, 2024), configured with a temperature of $1.0$ and default settings otherwise (as in the main paper for Claude 3 Haiku), produces $100$ candidate instructions from ten input-output examples per task; these are then embedded using BERT’s [CLS] token representation, and five representative instructions are selected via 5-medoid clustering. The validation split was used during optimization and the test split was used to assess unbiased performance. We perform standard sanitization of LLM outputs to be able to employ the loss function in Equation (1) based on the exact match as a scoring function. For GSM8K, we determine the prediction for the exact match loss function by selecting the last number contained in the LLM’s output. This approach aligns with the Chain-of-Thought prompting and the typical output behavior of LLMs for this task.
Table 5: Characteristics of tasks used in the experiments.
Task Setting $n_{\mathrm{train}}$ $n_{\mathrm{valid}}$ $n_{\mathrm{test}}$
AI2 ARC multiple choice question answering $1094$ $291$ $1144$
GSM8K grade school math questions $6154$ $1319$ $1319$
antonyms find antonym of word $2073$ $519$ $100$
larger animal select larger of two animals $2422$ $606$ $100$
negation negate a sentence $723$ $181$ $100$
object counting count number of objects $560$ $140$ $100$
orthography starts with output all words starting with a given letter $2400$ $600$ $100$
second word letter output the second letter of a word $2644$ $662$ $100$
sentiment sentiment analysis of movie rating $933$ $234$ $100$
word unscrambling build a word from scrambled letters $5627$ $1407$ $100$
@@ D.2 LLMs
We use Claude 3 Haiku (Anthropic, 2024), LLAMA3 8B Instruct (Llama Team, AI @ Meta, 2024), and Mistral 7B Instruct (Jiang et al., 2023) with default hyperparameters (Claude 3 Haiku: max tokens $=200$, temperature $=0.5$, top p $=1.0$, top $k=250$; LLAMA3 8B Instruct: max tokens $=512$, temperature $=0.5$, top $p=0.9$; Mistral 7B Instruct: max tokens $=512$, temperature $=0.5$, top p $=0.9$, top $k=50$). For GSM8K we increase max tokens for all LLMs to $1024$.
@@ D.3 Methods
We run all methods as described in Section 4. All full-fidelity BO methods are implemented within BoTorch (Balandat et al., 2020). We include HDBO (Hvarfner et al., 2024) to have a simple yet well-performing “high-dimensional” BO baseline. Hvarfner et al. (2024) recently challenged the general belief that vanilla BO does not perform well for high-dimensional functions by training a GP via MAP with priors over kernel and likelihood parameters adjusted to reflect the dimensionality of the problem which resulted in strong BO performance on high-dimensional functions. We run EASE as implemented in the official code base [^3] https://github.com/ZhaoxuanWu/EASE-Prompt-Optimization/blob/e3514de58bd682ebc5ea46fe890481f2b92e5589/experiments/LlamaForMLPRegression.py, MIPROv2 as implemented in DSPy [^4] https://github.com/stanfordnlp/dspy/blob/425b6f07d5cf0530f5a5566ad4f247b15aecb522/dspy/teleprompt/mipro_optimizer_v2.py and TRIPLE-SH and TRIPLE-GSE as implemented in the official code base [^5] https://github.com/ShenGroup/TRIPLE/blob/06264a97b4dd766c9a88afc24058627fac0f223d/src/bandit/contextual/gse.py. Regarding TRIPLE-GSE we noticed that Shi et al. (2024) did not provide detailed descriptions of the model used to predict expected prompt performance (in the main paper, they state that one can use a linear model or MLP). In their implementation, however, they use an ensemble of a Bayesian ridge regression model, a gradient boosting regression model, and an MLP where weights for the ensemble are determined based on each model’s $R^{2}$ performance on a separate validation set. We did not change this modeling approach when running TRIPLE-GSE. We implement our HbBoPs in GPyTorch (Gardner et al., 2018) and run it as described in Section 4. Moreover, to hedge against poor model-based proposals, we perform random interleaving as described in Falkner et al. (2018) for each proposal with a probability of $\rho=0.1$. Since a single SH or HB schedule may require less budget than the total pre-defined LLM call budget per task, we repeatedly run all methods that use SH or HB as multi-fidelity schedulers until their total LLM call usage reaches this pre-defined budget.
@@ E.1 Main Results
Here, we provide additional analyses of the main results reported in Section 5.1. To test whether HbBoPs outperforms all other methods with respect to validation and test error for different fractions of budget, we conduct a linear mixed effects model analysis. We model the unaggregated performance as a function involving random intercepts for each benchmark scenario (benchmark task and LLM combination). This approach is sensible as each method has been run repeatedly on the same benchmark scenario with different random seeds that affect, for example, initial designs. To test the global hypothesis that there is an effect of the method on performance, we test an intercept model against a model including an effect of the factor method. If we reject the null hypothesis, we proceed with a Tukey post hoc test (corrected for multiple testing) to test each method against HbBoPs. We test at the conservative $\alpha=0.01$ level.
For the validation error at a fraction of $1.00$, we reject the global null hypothesis of no effect of methods ($\chi^{2}(8)=672.95,p<1\mathrm{e}{-4}$). The pairwise results are:
@@
We conclude that HbBoPs outperforms all methods significantly with respect to final performance, except for TRIPLE-SH. While the effect is positive, i.e., HbBoPs improves over TRIPLE-SH, it is not strong enough to be considered statistically significant at the $\alpha=0.01$ level. For brevity, we do not include results for fractions of $0.25$ or $0.50$ here, where we observed similar results but HbBoPs more strongly outperforming the other methods.
Table 6: Normalized validation error of each method on each benchmark for Claude 3 Haiku. Averaged over repetitions. Standard errors in parentheses. For visual analysis, we highlight all methods that have a mean error that is less or equal to the mean error of the best method plus two times its standard error.
@@
Table 7: Normalized validation error of each method on each benchmark for LLAMA3 8B Instruct. Averaged over repetitions. Standard errors in parentheses. For visual analysis, we highlight all methods that have a mean error that is less or equal to the mean error of the best method plus two times its standard error.
@@
Table 8: Normalized validation error of each method on each benchmark for Mistral 7B Instruct. Averaged over repetitions. Standard errors in parentheses. For visual analysis, we highlight all methods that have a mean error that is less or equal to the mean error of the best method plus two times its standard error.
Benchmark AI2 ARC GSM8K antonyms larger animal negation object counting orthography starts with second word letter sentiment word unscrambling RS 0.329 (0.035) 0.160 (0.021) 0.020 (0.002) 0.043 (0.006) 0.006 (0.003) 0.040 (0.004) 0.164 (0.018) 0.171 (0.016) 0.098 (0.013) 0.224 (0.017) vanilla BO 0.359 (0.038) 0.166 (0.021) 0.017 (0.002) 0.040 (0.008) 0.008 (0.003) 0.033 (0.005) 0.117 (0.020) 0.136 (0.013) 0.085 (0.016) 0.161 (0.019) HDBO 0.217 (0.045) 0.105 (0.019) 0.013 (0.002) 0.021 (0.005) 0.009 (0.003) 0.028 (0.005) 0.120 (0.018) 0.085 (0.017) 0.042 (0.013) 0.108 (0.016) BOPCA 0.246 (0.044) 0.108 (0.019) 0.015 (0.002) 0.021 (0.006) 0.011 (0.004) 0.032 (0.006) 0.112 (0.021) 0.125 (0.019) 0.048 (0.013) 0.136 (0.021) EASE 0.384 (0.033) 0.133 (0.014) 0.022 (0.002) 0.053 (0.008) 0.000 (0.000) 0.035 (0.005) 0.147 (0.023) 0.183 (0.014) 0.110 (0.014) 0.095 (0.021) MIPROv2 0.323 (0.041) 0.122 (0.017) 0.013 (0.002) 0.032 (0.005) 0.011 (0.004) 0.032 (0.005) 0.125 (0.014) 0.158 (0.016) 0.068 (0.014) 0.163 (0.017) TRIPLE-SH 0.024 (0.011) 0.008 (0.003) 0.041 (0.005) 0.012 (0.003) 0.044 (0.008) 0.057 (0.006) 0.005 (0.004) 0.044 (0.012) 0.055 (0.010) 0.103 (0.018) TRIPLE-GSE 0.038 (0.016) 0.001 (0.000) 0.042 (0.005) 0.022 (0.006) 0.055 (0.010) 0.067 (0.010) 0.002 (0.000) 0.072 (0.015) 0.084 (0.017) 0.145 (0.022) HbBoPs 0.107 (0.036) 0.000 (0.000) 0.011 (0.001) 0.012 (0.003) 0.012 (0.005) 0.025 (0.005) 0.024 (0.010) 0.067 (0.013) 0.030 (0.011) 0.061 (0.014)
@@
We perform the same analysis for the test error at a fraction of $1.00$ and reject the global null hypothesis of no effect of methods ($\chi^{2}(8)=288.36,p<1\mathrm{e}{-4}$). The pairwise results are: